# Knowledge Distillation for Road Detection based on cross-model Semi-Supervised Learning

### 07 Feb 2024 v1

@@ Abstract
The advancement of knowledge distillation has played a crucial role in enabling the transfer of knowledge from larger teacher models to smaller and more efficient student models, and is particularly beneficial for online and resource-constrained applications. The effectiveness of the student model heavily relies on the quality of the distilled knowledge received from the teacher. Given the accessibility of unlabelled remote sensing data, semi-supervised learning has become a prevalent strategy for enhancing model performance. However, relying solely on semi-supervised learning with smaller models may be insufficient due to their limited capacity for feature extraction. This limitation restricts their ability to exploit training data. To address this issue, we propose an integrated approach that combines knowledge distillation and semi-supervised learning methods. This hybrid approach leverages the robust capabilities of large models to effectively utilise large unlabelled data whilst subsequently providing the small student model with rich and informative features for enhancement. The proposed semi-supervised learning-based knowledge distillation (SSLKD) approach demonstrates a notable improvement in the performance of the student model, in the application of road segmentation, surpassing the effectiveness of traditional semi-supervised learning methods.
@@ 1 Introduction
A significant challenge of (near-) real-time segmentation tasks like road surface detection is the trade-off between the accuracy of the segmentation and the speed at which it can be executed. More accurate methods generally require more model parameters and higher running time in evaluation. Whilst employing small models often leads to quicker task execution, they may lack the capacity to efficiently grasp and represent knowledge concisely when provided with comparable computational resources and data as larger models. To solve this problem, knowledge distillation is used to transfer knowledge from large models to smaller ones without loss of validity [10]. However, nowadays, the majority of knowledge distillation studies in the field of remote sensing primarily focus on employing labeled data for training both teacher and student models, while not sufficiently investigating the use of unlabeled data to enhance the performance of student models.
@@
Specifically, the contributions of this work are as follows: (1) We introduce a knowledge distillation process and a framework comprising two distinct teacher models to enhance the effectiveness of a lightweight and fast student model designed for road segmentation. Essentially, the framework undertakes an extensive knowledge distillation across various levels, encompassing features, probabilities, and labels. Moreover, this distilled knowledge can be consistently applied to diverse segmentation tasks. (2) We employ a combination of semi-supervised learning and knowledge distillation methodologies. This strategy involves leveraging a significant volume of unlabelled data to enhance the capabilities of the lightweight student model. Simultaneously, the larger teacher models contribute comprehensive and informative guidance for the process of knowledge distillation.
@@ 2 Related Work
Fig. 2: Framework of Cross-model Supervision
Fig. 3: Framework of the proposed semi-supervised learning based knowledge distillation (SSLKD)
@@
$\mathcal{L}_{sup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k},y_{k}\right)\right),$ (1)
@@
Following this, to further improve the two teacher models’ representation to provide better features and pseudo labels for knowledge distillation, we implemented cross-model supervision to train teacher models by using a large amount of unlabelled data. This is due to the fact that different networks typically possess complementary features and predictions, and they are regarded as teachers for one another. The framework of cross-model supervision is shown in Figure 2 where Teacher #1 refers to a DeepLabV3+ [18] architecture with a Resnet-101 backbone, and Teacher #2 represents a SegNet [16] with a VGG-16 backbone. Specifically, the predictions made by both teachers are employed to create pseudo labels, which are then utilised in training the other network through the cross-entropy loss function; however, these labels are not employed in the process of self-training.
@@
$\mathcal{L}_{dis}^{f}=\frac{1}{C\times W\times H}\sum_{k=1}^{C\times W\times H}\left(|f_{i}-\hat{f}_{i}|\right),$ (2)
@@
Table 1: Performance Comparison of Different Methods for the RoadNet Dataset.
@@ 4 EXPERIMENTS AND RESULTS
The performance of teacher and student models trained by supervised learning are shown in the top three lines of Table 1. Both of the two teacher models show superior performance compared to the student model as expected. Notably, the intersection over union (IoU) for Teacher #2 is approximately 4% higher than that of the student model. However, the values of GFLOPs for the teacher models exceed that of the student model. There is a significant enhancement in their performance after implementing cross-model supervision with unlabelled data for both teacher models. For instance, Teacher #1 exhibits a 3.64% improvement in IoU (line 4 of Table 1).

### 25 Mar 2024 v2

@@ Abstract
The advancement of knowledge distillation has played a crucial role in enabling the transfer of knowledge from larger teacher models to smaller and more efficient student models, and is particularly beneficial for online and resource-constrained applications. The effectiveness of the student model heavily relies on the quality of the distilled knowledge received from the teacher. Given the accessibility of unlabelled remote sensing data, semi-supervised learning has become a prevalent strategy for enhancing model performance. However, relying solely on semi-supervised learning with smaller models may be insufficient due to their limited capacity for feature extraction. This limitation restricts their ability to exploit training data. To address this issue, we propose an integrated approach that combines knowledge distillation and semi-supervised learning methods. This hybrid approach leverages the robust capabilities of large models to effectively utilise large unlabelled data whilst subsequently providing the small student model with rich and informative features for enhancement. The proposed semi-supervised learning-based knowledge distillation (SSLKD) approach demonstrates a notable improvement in the performance of the student model, in the application of road segmentation surpassing the effectiveness of traditional semi-supervised learning methods.
@@ 1 Introduction
A significant challenge of (near-) real-time segmentation tasks like road surface detection is the trade-off between the accuracy of the segmentation and the speed at which it can be executed. More accurate methods generally require more model parameters and higher running time for inference. Whilst employing small models often leads to quicker task execution, they may lack the capacity to efficiently grasp and represent knowledge when provided with comparable computational resources and data as larger models. To solve this problem, knowledge distillation is used to transfer knowledge from large models to smaller ones without loss of validity [10]. However, nowadays, the majority of knowledge distillation studies in the field of remote sensing primarily focus on employing labeled data for training both teacher and student models, while not sufficiently investigating the use of unlabeled data to enhance the performance of student models.
@@
Specifically, the contributions of this work are as follows: (1) We introduce a knowledge distillation process and a framework comprising two distinct teacher models to enhance the effectiveness of a lightweight and fast student model designed for road segmentation. Essentially, the framework undertakes extensive knowledge distillation across various levels, encompassing features, probabilities, and labels. Moreover, this distilled knowledge can be consistently applied to diverse segmentation tasks. (2) We employ a combination of semi-supervised learning and knowledge distillation methodologies. This strategy involves leveraging a significant volume of unlabelled data to enhance the capabilities of the lightweight student model. Simultaneously, the larger teacher models contribute comprehensive and informative guidance for the process of knowledge distillation.
@@ 2 Related Work
Fig. 2: Framework of Cross-model Supervision.
Fig. 3: Framework of the proposed semi-supervised learning based knowledge distillation (SSLKD).
@@
$\mathcal{L}_{sup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k},y_{k}\right),$ (1)
@@
Following this, to further improve the two teacher models’ representation to provide better features and pseudo labels for knowledge distillation, we implemented cross-model supervision to train the teacher models by using a large amount of unlabelled data. This is due to the fact that different networks typically possess complementary features and predictions, and they are regarded as teachers for one another. The framework of cross-model supervision is shown in Figure 2 where Teacher #1 refers to a DeepLabV3+ [18] architecture with a Resnet-101 backbone, and Teacher #2 represents a SegNet [16] with a VGG-16 backbone. Specifically, the predictions made by both teachers are employed to create pseudo labels, which are then utilised in training the other network through the cross-entropy loss function; however, these labels are not employed in the process of self-training.
@@
$\mathcal{L}_{dis}^{f}=\frac{1}{C\times W\times H}\sum_{k=1}^{C\times W\times H}|f_{i}-\hat{f}_{i}|,$ (2)
@@
Table 1: Comparing the performance of individual student and teacher models, along with students trained using the proposed method and other techniques, on the RoadNet dataset.
@@ 4 EXPERIMENTS AND RESULTS
The performance of teacher and student models trained separately by supervised learning are shown in the top three lines of Table 1. Both of the two teacher models show superior performance compared to the student model as expected. Notably, the intersection over union (IoU) for Teacher #2 is approximately 4% higher than that of the student model. However, the values of GFLOPs for the teacher models exceed that of the student model. There is a significant enhancement in their performance after implementing cross-model supervision with unlabelled data for both teacher models. For instance, Teacher #1 exhibits a 3.64% improvement in IoU (line 4 of Table 1).